OpenStudy (anonymous):

A painter bought 8 gallons of paint, just enough to cover two rooms. One room is twice the size of the other and requires three coats of paint. The smaller room requires only two coats. Then the client decided not to paint the smaller room after all. Let X be the number of gallons it takes for one coat of paint in the large room and Y the number for one coat of paint in the small room. How many gallons of paint will be left? @agent0smith

OpenStudy (anonymous):

x + y = 8 x = 2y 2y + y = 8 3y = 8 y = 8/3 = 2 2/3 x = 2y x = 2(8/3) x = 16/3 = 5 1/3 check... x + y = 8 8/3 + 16/3 = 8 24/3 = 8 8 = 8 (correct) large room uses 5 1/3 gallons smaller room uses 2 2/3 gallons However, the smaller room was not painted, therefore, 2 2/3 gallons were left over. I am not 100% sure on this, so wait until you can get a second opinion

OpenStudy (anonymous):

The correct answer was 2.

OpenStudy (anonymous):

crap.....I wonder how they got that answer....do you know ?